# EARN Fairness: Explaining, Asking, Reviewing and Negotiating Artificial Intelligence Fairness Metrics Among Stakeholders

### v1

@@ 3.2.3. UI components
Static Fairness Explanation and Exploration: This component offers interactive visualizations of fairness metrics and assessment results of AI fairness (Figure 1 C), including three panels: ”Group Fairness Overview”, ”Subgroup Fairness Overview”, and ”Individual Fairness Overview”. Each panel works in a similar way, initially presenting a comprehensive overview displaying all metrics’ results within this fairness category. In light of prior work (Nakao et al., 2023), users can adjust the fairness metric threshold via a slider, dynamically altering a green rectangle that denotes the fairness zone. Metrics within the green area appear as green dots, indicating the model is fair at the current threshold. Metrics outside the area are red dots, signaling the model is unfair. For group and subgroup fairness, users can switch between different protected features or their combinations to assess fairness. Users can decide on their threshold settings and select protected features or combinations of interest in these panels.
Next, by clicking the ”VIEW EXPLANATIONS” button within each panel, users can access detailed explanations for metrics under each fairness category and select their top three preferred metrics. These explanations use text, charts, and images to break down fairness metrics at the individual case level, reducing the need for a mathematical or statistical background. Specifically:
@@ 3.2.4. Comparison with previous fairness UIs
Leveraging Nakao et al.’s work (Nakao et al., 2023) which provided user requirements of fairness investigation tools in a loan scenario, Table 1 summarizes a comparison between our UI, FEE, and two other tools, FET (Cheng et al., 2021) and FairHil (Nakao et al., 2023). (We extracted AI fairness exploration requirements from loan officers’ perspective, as we believe this group more closely aligns with our target audience, who are stakeholders without AI knowledge). Our tool, FEE, differs by supporting custom threshold inputs and dynamically explaining when metrics indicate fairness or unfairness. It also enhances the transparency of AI fairness by providing detailed model parameters and instance-level fairness metric explanations.
@@ 4.1. Data and Model
We employed a logistic regression classifier. Logistic regression’s simplicity and interpretability make it ideal for improving model understanding, particularly for non-experts (Barredo Arrieta et al., 2020). Previous research has also effectively employed logistic regression to study fairness in this dataset (Hort et al., 2023). We used all 1000 instances, training this model with 5-fold cross-validation to ensure sufficient test instances for subgroup fairness analysis. To balance model accuracy and fairness, our model achieves a 0.76 accuracy on 200 test instances.
@@ 4.3. Procedure
Finally, we followed the framework to conduct team sessions for each team to complete Review and Negotiate. Members were provided with hard copies outlining negotiation objectives and engaged in free-form negotiation without interference from researchers. During the negotiation process, members were free to use our interactive system. The negotiation results should be unanimously approved by all team members. Researchers aimed to ensure a harmonious negotiation atmosphere and offered alternative collective decision-making methods when needed, such as the Borda voting method due to its simplicity and robustness (Lee et al., 2019b; Kahng et al., 2019).
@@ 4.4. Data Collection and Analysis
We collected participants’ Top 3 Metric Ranking Lists and fairness thresholds, along with audio recordings and the researcher’s observation notes. We used descriptive statistics to analyze the Top 3 Metric Ranking Lists, reporting the raw frequencies for each metric within the top 3 ranks, and to calculate the mean values and variances for fairness thresholds. Further, we transcribed approximately 20 hours of recordings from all 18 participants. We conducted a thematic analysis (Braun and Clarke, 2006) to generate insights into participants’ preferences on fairness metrics and thresholds, the reasons behind them, and user feedback. Our research team met weekly to iterate over the themes and coding, reviewing and validating them. After 6 rounds of iteration, we finalized the final codebook (Appendix A.3).
@@ 5. Results
5.1. What are stakeholders’ personal preferences on AI fairness metrics, and why? (RQ1)
@@ 5.1.1. Preferences for Fairness Metrics
Table 2 shows all the 18 participants’ preferences for fairness metrics and fairness thresholds. We first investigated preferred fairness metric categories. Individual fairness emerged as the most favored fairness metric category, chosen by 9 participants in their Top-1 metric, closely followed by subgroup fairness, chosen by 7 participants. In contrast, despite group fairness being widely favored in industry and academia for its ease of implementation (Binns, 2020), it only garnered support from 2 participants in this study.
@@
Then we delved into specific fairness metrics. Figure 5 shows the ranking distribution of metrics chosen by participants. We observed that preferences diverged among the participants’ Top-1 choices. Conditional Statistical Parity was the most preferred, chosen by 7 out of 18 participants, which indicates the importance stakeholders place on non-protected features (job, savings, employment, and credit history in our credit rating scenario) in fairness assessment. It was followed by Counterfactual Fairness (6 participants), and Consistency (5 participants), indicating preferences for fairness at the individual level. It is noteworthy that Demographic Parity, the most widely used metric in existing AI fairness research (Hort et al., 2023; Verma and Rubin, 2018; Mehrabi et al., 2021), was not chosen by any participant as their Top-1 choice. Equal Opportunity and Equalized Odds, which are commonly used fairness metrics, were each selected by only one participant.
@@
Overall, while there were subgroups of participants that seemed very similar in their preferences, there was a lack of consensus. For example, P10, P11, P16 and separately P6, P14 had identical selections of preferred metrics in all the top three ranks but did not agree with each other. Participants did not agree on fairness metric categories, let alone fairness metrics within them, making how to measure and audit fairness problematic.
Therefore, a strategy for choosing fairness metrics could instead leverage stakeholders’ ranked preferences, especially as we observed that each metric was included in the top three rankings. To investigate what metric was most highly ranked overall, we decided to weight these rankings to investigate overall patterns of preference, with weights of 3, 2, and 1 for the Top-1, Top-2, and Top-3 choices, respectively. We found that Conditional Statistical Parity (36 points) stands out far ahead of the other metrics in terms of ranked preference, with Consistency (23 points) and Counterfactual Fairness (21 points) following behind. Other metrics, such as Equalized Odds (13 points), Equal Opportunity (12 points), Predictive Equality (5 points), Outcome Test (5 points) and Demographic Parity (2 points) were ranked a lot lower.
Lastly, we also gathered user preferences for protected features and feature combinations when applying group and subgroup fairness. The majority of participants identified age as their primary concern (8 participants), followed by gender, which was selected by seven participants. Interestingly, only three participants, all of whom were foreign residents, selected foreign worker status as their primary concern.
Note that 14 out of 18 participants preferred subgroup fairness over group fairness. Consequently, the combined consideration of age and gender was identified by a majority of participants (7 participants). Three participants advocated for an even broader inclusivity, urging the consideration of age, gender, and foreign worker status at the same time. Less commonly selected combinations were gender and foreign worker status (2 participants), or age and foreign worker status (1 participant).
Main finding 1: In credit prediction scenarios, lay stakeholders favored the individual-level fairness metric category, and Conditional Statistical Parity stood out as the preferred metric. Overall, lay stakeholders’ preferences for fairness metrics diverged significantly from the current popular metrics in industry and academia. There was strong support for focusing on age and gender attributes and their intersection.
@@
Table 2. Participants’ Preferences for AI Fairness Metrics and Fairness Thresholds. Top-1, Top-2, and Top-3 represent participants’ top three preferred metrics. Group Fairness vs. Subgroup Fairness indicates whether participants prefer the selected applying group-level fairness metric(s) to group fairness or subgroup fairness. Group and subgroup fairness thresholds deem the model fair if differences are below these values. Individual fairness thresholds deem the model fair if results exceed these values.
@@ 5.1.2. Reasons for Metric Preferences and Concerns
We analyzed participants’ think-aloud data from individual sessions using a codebook (Appendix A.3.1) and highlight the code in this section in bold text. We expected these reasons to be reprised and discussed during group negotiations as arguments for a metric consensus.
@@
Main finding 2: Participants preferred individual fairness for acknowledging individual differences overlooked by group-level metrics, and favored subgroup fairness over group fairness for minimizing differences within groups. Conditional Statistical Parity stood out for valuing non-protected features that are influential in AI fairness decisions, promoting evaluations based on capabilities, despite proximity concerns between non-protected and protected features. Consistency was notable for allowing comparisons among similar people, promoting fairness and inclusivity, though participants struggled with measuring similarity objectively. Counterfactual Fairness was praised for concentrating on individual traits and objective quantification, though doubts were raised about its real-world efficacy.
5.1.3. Fairness Thresholds and Reasons
We show how participants set these thresholds in Table 2. We also analyzed the reasons for choosing these thresholds during individual sessions; our codebook can be found in Appendix A.3.2.
We observed that there is a wide variability of thresholds, from 0% to 30% in group fairness, 0% to 30% in subgroup fairness, and 80% to 100% in individual fairness. Further, no two participants set the exact same thresholds for the three fairness categories. In group and subgroup fairness settings, only two participants chose a 0% threshold, while four opted for a 100% threshold in individual fairness. One approach for agreeing on these could be to average them. The mean for group fairness was 9.28% (sd=7.51%), subgroup fairness was 13.00% (sd=9.96%), and individual fairness was 92.44% (sd=8.53%). We found that participants imposed the strictest thresholds for individual fairness; within the context of group-level fairness, we observed that thresholds for subgroup fairness were set more leniently than group fairness  [^3] Based on the metrics’ calculation, for individual fairness, unfairness is 100% minus the threshold, so a 92.44% threshold means 7.56% unfairness is acceptable. For group and subgroup fairness, the thresholds directly represent allowable (sub)group differences, which are 9.28% and 13.00%, respectively.. We can note that on average participants set thresholds much higher than expected by law.
Most participants were willing to leave some room for unfairness, as it was deemed impossible to achieve in practice. As P14 remarked, ”We don’t need to be too absolute, … Experts or AI are not 100% accurate, that absolute perfection in fairness and accuracy is unattainable”. Only a small number of participants pursued perfect fairness, reflecting a zero-tolerance attitude towards bias. One reason for the higher thresholds for subgroup fairness was that more features were being considered, thereby allowing for more leeway. As P10 stated, ”Since you’re taking more factors into account when assessing fairness, like more individual factors, you can maybe increase this a bit more, like […] more leeway”. Strangely, participants did not extend this reasoning to individual fairness, even though even more features are considered in this fairness metric. This reflects higher fairness expectations or lower unfairness tolerance due to the majority of participants preferring the individual fairness category.
Main finding 3: Participants set stringent standards for individual fairness, reflecting their higher expectations. However, they allowed more flexibility in subgroup fairness due to the consideration of additional features compared to group fairness.
@@ 5.2. How do stakeholders negotiate to resolve personal differences in metric preferences? (RQ2)
Because stakeholders have diverse views and preferences on metrics and fairness categories, reflecting these views in ”fairer” AI models remains challenging. In this section, we investigate how team members negotiate to resolve differences in metric preferences, identifying three distinct consensus-reaching styles without researcher input. We will detail the negotiation strategies shaped by team decision-making and leadership, and the consensus each team ultimately reached.
@@
In Team 1, participants adopted mainly a majority-based voting approach to resolve their personal metric preferences. In this approach, they narrowed down options based on the frequency of preferred metrics. Team members agreed to follow the majority and abandon any minority preferences. They started out by choosing to prioritize subgroup fairness over group fairness, following a majority vote suggested by P11. Next, they eliminated the three least chosen metrics, as suggested by P8. Further, as proposed by P4, they narrowed down the selection of the specific metric to employ to the Top-1 choice, either Conditional Statistical Parity or Counterfactual Fairness. At this point, P7, P11, and P12 favored Counterfactual Fairness due to their personal Top-1 preference for individual-level fairness, while P4, P5, and P8 supported Conditional Statistical Parity, driven by their Top-1 preference for group-level fairness. This then led to in-depth discussions between team members as to the reasons behind their Top-1 choices, resurfacing what they had talked about during their earlier one-on-one sessions (Section 5.1.2). We also found that the process of discussion at this stage stimulated the exploration of the core advantages and disadvantages of fairness and ways to measure it, thus deepening participants’ understanding of the metrics. For example, P7 supported Counterfactual Fairness and pointed out the drawbacks of group fairness: ”The thing is the subgroups just level up other groups, right? So if you and I have the same credit score and we apply for a credit, […], as I said, I don’t get the credit but you get it because the metrics is 50% of all males good credit, only 30% of the females good credits. So we have to give more credits to females. So we have exactly the same income, but [you]’ll be prioritized because you belong to a group […] [What] group and subgroup [metrics] generally do is to try from a government perspective, equalized. But the question is, is it fair or not?”. On the other hand, P8, a supporter of Conditional Statistical Parity, expressed concerns about individual fairness, stressing the importance of group fairness in maintaining equity between protected and non-protected groups at the societal level: ”It’s maybe unfair […] [Unless] I’m selfish, I will go with individual fairness”. Meanwhile, P11 emphasized the decision not to use Consistency due to the complexity of similarity calculations, explaining, ’Instead of neighbor because you have to explain again what this neighbor means, it is that you have the same [information]. But if the counterfactual fairness, it’s just like male [or] female […] or something like that’. This nuance wasn’t fully realized during P11’s individual session, yet group discussions further deepened the understanding of this metric. Therefore, after this discussion, the team agreed on a ”hybrid” solution: using Conditional Statistical Parity with a focus on subgroup fairness first, and then using Counterfactual Fairness as a backup for individual fairness. As P8 suggested: ”[Use] Conditional Statistical Parity first and then I when I see a complaint then I would go with the Counterfactual Fairness to decide to, to judge him a second time”.
@@
Team 2, similar to Team 1, also tried to discuss the reasons for their choices and then vote on them, emphasizing collective agreement to reconcile their diverging preferences. However, this discussion led to entrenched positions and inhibited negotiation. For example, initially, P1 suggested that each team member share and justify their top 3 preferred metrics in turn. Then, they checked if any team member had changed their mind after hearing the others. However, all members stuck to their initial personal choices. Then, as suggested by P17, they tried to overcome that deadlock by voting on the team’s top 3 metric list. However, the team only accepted unanimous decisions and therefore this approach failed. As a final way out, the team attempted a compromise solution. P17 suggested: ”I like the idea of grouping with the Conditional Statistical Parity and then choosing an individual fairness afterward. And within same condition groups that we want to make sure each individual can be treated fairly”. All members, except P15, favored individual fairness with Counterfactual Fairness, while P15 suggested a combination of Counterfactual Fairness and Consistency. For comprehensiveness, they chose to combine both individual fairness metrics. Team 2 further prioritized subgroup fairness for Conditional Statistical Parity using a majority-based approach. Therefore, similar to Team 1, Team 2 also agreed on a ”hybrid” solution: using Conditional Statistical Parity with a focus on subgroup fairness first, and then using Counterfactual Fairness and Consistency as backups for individual fairness.
@@
In Team 3, reasons for and against fairness metrics were also discussed but they used a debate style, encouraging the presentation of arguments for and against certain fairness metrics. In the early negotiation stages, some polarization emerged similar to Team 1, but was successfully overcome through argumentation. For example, P10, P13, and P14 supported Counterfactual Fairness, while P9 and P18 favored Conditional Statistical Parity. In this team, we observed that certain team members dominated the discussions, leading to significant variations in the levels of participation among the group members. P10 presented arguments for Counterfactual Fairness while P18 supported Conditional Statistical Parity, leading two sides. The discussion primarily revolved around P18’s focus on group-level legitimate factors for fairness versus P10’s approach of directly excluding the impact of protected attributes for AI decision-making fairness. P10 proposed fairness as follows: ”It’s like oh no matter the value of age, gender, foreign worker is changed or not, AI prediction should not be changed […] You eliminate the bias first and then you think about employment [and other individual qualifications]”. After this perspective was accepted, Team 3, different from Team 1 and Team 2, advocated for using Counterfactual Fairness for individual fairness first, and then using Conditional Statistical Parity with a focus on subgroup fairness. Additionally, there was unanimous agreement on the importance of including at least age and gender to secure subgroup fairness in the Conditional Statistical Parity application.
@@
Main finding 4: Teams used different strategies to come to a consensus. Voting was frequently employed but there were differences in how to decide the outcomes of voting. Reasons for or against metrics were discussed in all teams but this can also lead to entrenchments of opinions if not carefully handled. All teams chose to overcome conflict by combining metrics in some way, ending up with compromise solutions. The consensus showed that all teams considered both subgroup fairness and individual fairness, but with different priorities.

### v2

@@ 3.2.3. UI components
Static Fairness Explanation and Exploration: This component offers interactive visualizations of fairness metrics and assessment results of AI fairness (Figure 1 C), including three panels: ”Group Fairness Overview”, ”Subgroup Fairness Overview”, and ”Individual Fairness Overview”. Each panel works in a similar way, initially presenting a comprehensive overview displaying all metrics’ results within this fairness category. For group and subgroup fairness, users can select and switch between different protected features or their combinations to assess fairness. In light of prior work (Nakao et al., 2023), users can also adjust the fairness threshold for each category by dragging a slider, which dynamically alters a green rectangular area representing the ”fairness zone”. Metrics inside this green area, appearing as green dots, indicate the model is fair at the current threshold. Metrics outside the area are red dots, signaling the model is unfair. During this process, users can intuitively visualize AI fairness assessment results for the current metric, compare fairness outcomes across different metrics and categories to set fairness thresholds, or consult detailed explanations further to refine their threshold settings.
By clicking the ”VIEW EXPLANATIONS” button within each panel, users can next access detailed explanations for metrics under each fairness category and select their top three preferred metrics. These explanations use text, charts, and images to break down fairness metrics at the individual case level, reducing the need for a mathematical or statistical background. Specifically:
@@ 3.2.4. Comparison with previous fairness UIs
Leveraging Nakao et al.’s work (Nakao et al., 2023) which provided user requirements of fairness investigation tools in a loan scenario  [^1] We extracted AI fairness exploration requirements from loan officers’ perspective, as we believe this group more closely aligns with our target audience, who are stakeholders without AI knowledge, Table 1 summarizes a comparison between our UI, FEE, and two other tools, FET (Cheng et al., 2021) and FairHIL (Nakao et al., 2023). Our tool, FEE, differs by supporting custom threshold inputs and dynamically explaining when metrics indicate fairness or unfairness. It also enhances the transparency of AI fairness by providing detailed model parameters and instance-level fairness metric explanations.
@@ 4.1. Data and Model
We employed a logistic regression classifier. Logistic regression’s simplicity and interpretability make it ideal for improving model understanding, particularly for non-experts (Barredo Arrieta et al., 2020). Previous research has also effectively employed logistic regression to study fairness in this dataset (Hort et al., 2024). We used all 1000 instances, training this model with 5-fold cross-validation to ensure sufficient test instances for subgroup fairness analysis. To balance model accuracy and fairness, our model achieves a 0.76 accuracy on 200 test instances.
@@ 4.3. Procedure
Finally, we followed the framework to conduct team sessions for each team to complete Review and Negotiate. Members were provided with hard copies outlining their task and engaged in free-form negotiation without interference from researchers. Guidance was provided that the negotiation results should be unanimously approved by all team members. During the negotiation process, members were free to use FEE. Researchers aimed to ensure a harmonious negotiation atmosphere and offered alternative collective decision-making methods when needed, such as the Borda voting method due to its simplicity and robustness (Lee et al., 2019b; Kahng et al., 2019).
@@ 4.4. Data Collection and Analysis
We collected participants’ top 3 metric ranking lists and fairness thresholds, along with audio recordings and the researcher’s observation notes. We used descriptive statistics to analyze the top 3 metric ranking lists, reporting the raw frequencies for each metric within the top 3 ranks, and to calculate the mean values and variances for fairness thresholds. Further, we transcribed approximately 20 hours of recordings from all 18 participants. We conducted a thematic analysis (Braun and Clarke, 2006) to generate insights into participants’ diverse preferences on fairness metrics and thresholds, the reasons behind them, and user feedback on FEE design. Our research team met weekly to iterate over the themes and coding, reviewing and validating them. After 6 rounds of iteration, we finalized the final codebook (Appendix A.3).
@@ 5. Results
5.1. What are stakeholders’ personal preferences on AI fairness metrics that require negotiation? (RQ1)
@@ 5.1.1. Preferences for Fairness Metrics
Table 2 shows all the 18 participants’ preferences for fairness metrics and fairness thresholds. We first investigated preferred fairness metric categories that demonstrate differences to be negotiated. Individual fairness emerged as the slightly more favored fairness metric category, chosen by 9 participants in their Top-1 metric, closely followed by subgroup fairness, chosen by 7 participants. In contrast, despite group fairness being widely favored in industry and academia for its ease of implementation (Binns, 2020), it only garnered support from 2 participants in this study.
@@
Then we delved into specific fairness metrics. Figure 5 shows the ranking distribution of metrics chosen by participants. We observed that preferences diverged among the participants’ Top-1 choices. Conditional Statistical Parity was the most preferred, chosen by 7 out of 18 participants, which indicates the importance stakeholders might place on non-protected features (job, savings, employment, and credit history in our credit rating scenario) in fairness assessment. It was followed by Counterfactual Fairness (6 participants), and Consistency (5 participants), indicating preferences for fairness at the individual level. It is noteworthy that Demographic Parity, the most widely used metric in existing AI fairness research (Hort et al., 2024; Verma and Rubin, 2018; Mehrabi et al., 2021), was not chosen by any participant as their Top-1 choice. Equal Opportunity and Equalized Odds, which are commonly used fairness metrics, were each selected by only one participant.
We also observed that there is a wide variability of thresholds in Table 2, driven by a range of reasons (Appendix A.3.1). No two participants set exactly the same thresholds for all three fairness categories. In group and subgroup fairness settings, only two participants chose a 0% threshold, while four opted for a 100% threshold in individual fairness. We can note that on average participants set thresholds relatively higher than expected by law and by most practices (Gupta et al., 2021), with a group fairness mean of 9.28% (SD=7.51%), a subgroup fairness of 13.00% (SD=9.96%), and an individual fairness of 92.44% (SD=8.53%).
Most participants were willing to leave some room for unfairness, as it was deemed impossible to achieve in practice. As P14 remarked, ”We don’t need to be too absolute, … Experts or AI are not 100% accurate, that absolute perfection in fairness and accuracy is unattainable”. However, a small number of participants pursued perfect fairness, reflecting a zero-tolerance attitude towards bias. Another interesting observation is that participants tended to set higher thresholds for subgroup fairness, as more features were being considered compared to group fairness, thereby allowing for more leeway. As P10 stated, ”Since you’re taking more factors into account when assessing fairness, like more individual factors, you can maybe increase this a bit more, like […] more leeway”.
@@
Overall, while there were subgroups of participants that seemed very similar in their preferences, there was a lack of consensus. For example, P10, P11, P16 and separately P6, P14 had identical selections of preferred metrics in all the top three ranks but did not agree with each other. Participants did not agree on fairness metric categories, let alone fairness metrics within them, making reaching consensus on fairness problematic.
Main finding 1: Lay stakeholders’ preferences for fairness metrics appeared to differ from the current popular metrics in industry and academia, and more importantly there was no consensus for one metric or metric category.
@@
Table 2. Participants’ Preferences for AI Fairness Metrics and Fairness Thresholds. Top-1, Top-2, and Top-3 represent participants’ top three preferred metrics. Group Fairness vs. Subgroup Fairness indicates whether participants prefer applying the selected group-level fairness metric(s) to group fairness or subgroup fairness. Group and subgroup fairness thresholds deem the model fair if differences are below these values. Individual fairness thresholds deem the model fair if results exceed these values.
@@ 5.1.2. Reasons for Metric Preferences and Concerns
We analyzed participants’ think-aloud data from individual sessions using a codebook (Appendix A.3.2) and highlighted the code in this section in bold text. We expected these reasons to be reprised and discussed during group negotiations as arguments for a metric consensus, as detailed in Section 5.2.
@@
Main finding 2: There were numerous justifications for choosing particular metrics among participants. Participants who preferred individual fairness often pointed to individual differences overlooked by group-level metrics. Participants also placed importance on non-protected features that influence AI fairness decisions and emphasized evaluations based on capabilities, which motivated them to choose Conditional Statistical Parity. Participants who chose Consistency noted that it allowed comparisons among similar people, promoting fairness and inclusivity. Counterfactual Fairness was praised for concentrating on individual traits and objective quantification.
@@ 5.2. How do stakeholders negotiate to resolve personal differences in metric preferences? (RQ2)
Because stakeholders have diverse views and preferences on metrics and fairness categories, reflecting these views in ”fairer” AI models remains challenging. In this section, we investigate how team members negotiated to resolve differences in metric preferences after using FEE to review and build a shared understanding of each selected metric. We identified three distinct negotiation strategies for consensus-reaching, achieved without researcher input. We will detail the negotiation strategies shaped by team decision-making and leadership, and the consensus each team ultimately reached.
@@
In Team 1, participants adopted mainly a majority-based voting approach to resolve their personal metric preferences. In this approach, they narrowed down options based on the frequency of preferred metrics. Team members agreed to follow the majority and abandon any minority preferences. They started out by choosing to prioritize subgroup fairness over group fairness, following a majority vote suggested by P11. Next, they eliminated the three least chosen metrics, as suggested by P8. Further, as proposed by P4, they narrowed down the selection of the specific metric to employ to the Top-1 choice, either Conditional Statistical Parity or Counterfactual Fairness. At this point, P7, P11, and P12 favored Counterfactual Fairness due to their personal Top-1 preference for individual-level fairness, while P4, P5, and P8 supported Conditional Statistical Parity, driven by their Top-1 preference for group-level fairness. This then led to in-depth discussions between team members as to the reasons behind their Top-1 choices. During the discussion, team members consulted FEE for insights on Conditional Statistical Parity and Counterfactual Fairness, resurfacing and sharing what they had talked about during their earlier individual sessions (Section 5.1.2). We also found that the process of discussion at this stage stimulated the exploration of the core advantages and disadvantages of fairness and ways to measure it, thus deepening participants’ understanding of the metrics. For example, P7 supported Counterfactual Fairness and pointed out the drawbacks of group fairness: ”The thing is the subgroups just level up other groups, right? So if you and I have the same credit score and we apply for a credit, […], as I said, I don’t get the credit but you get it because the metrics is 50% of all males good credit, only 30% of the females good credits. So we have to give more credits to females. So we have exactly the same income, but [you]’ll be prioritized because you belong to a group […] [What] group and subgroup [metrics] generally do is to try from a government perspective, equalized. But the question is, is it fair or not?”. On the other hand, P8, a supporter of Conditional Statistical Parity, expressed concerns about individual fairness, stressing the importance of group fairness in maintaining equity between protected and non-protected groups at the societal level: ”It’s maybe unfair […] [Unless] I’m selfish, I will go with individual fairness”. Meanwhile, P11 emphasized the decision not to use Consistency due to the complexity of similarity calculations, explaining, ’Instead of neighbor because you have to explain again what this neighbor means, it is that you have the same [information]. But if the counterfactual fairness, it’s just like male [or] female […] or something like that’. This nuance was not fully realized during P11’s individual session, yet group discussions further deepened the understanding of this metric. Therefore, after this discussion, the team agreed on a ”hybrid” solution: using Conditional Statistical Parity with a focus on subgroup fairness first, and then using Counterfactual Fairness as a backup for individual fairness. As P8 suggested: ”[Use] Conditional Statistical Parity first and then I when I see a complaint then I would go with the Counterfactual Fairness to decide to, to judge him a second time”.
@@
Team 2, similar to Team 1, also tried to discuss the reasons for their choices and then vote on them, emphasizing collective agreement to reconcile their diverging preferences. However, this discussion led to entrenched positions and inhibited negotiation. For example, initially, P1 suggested that each team member share and justify their top 3 preferred metrics in turn. In this process, each one used FEE to demonstrate their metric preferences. Then, they checked if any team member had changed their mind after hearing the others. However, all members stuck to their initial personal choices. Then, as suggested by P17, they tried to overcome that deadlock by voting on the team’s top 3 metric list. However, the team only accepted unanimous decisions and therefore this approach failed. As a final way out, the team attempted a compromise solution. P17 suggested: ”I like the idea of grouping with the Conditional Statistical Parity and then choosing an individual fairness afterward. And within same condition groups that we want to make sure each individual can be treated fairly”. All members, except P15, favored individual fairness with Counterfactual Fairness, while P15 suggested a combination of Counterfactual Fairness and Consistency. For comprehensiveness, they chose to combine both individual fairness metrics. Team 2 further prioritized subgroup fairness for Conditional Statistical Parity using a majority-based approach. Therefore, similar to Team 1, Team 2 also agreed on a ”hybrid” solution: using Conditional Statistical Parity with a focus on subgroup fairness first, and then using Counterfactual Fairness and Consistency as backups for individual fairness.
@@
In Team 3, reasons for and against fairness metrics were also discussed but they used a debate style, encouraging the presentation of arguments for and against certain fairness metrics. In the early negotiation stages, some polarization emerged similar to Team 1, but was successfully overcome through argumentation. For example, P10, P13, and P14 supported Counterfactual Fairness, while P9 and P18 favored Conditional Statistical Parity. In this team, we observed that certain team members dominated the discussions, leading to significant variations in the levels of participation among the group members. P10 presented arguments for Counterfactual Fairness while P18 supported Conditional Statistical Parity, leading two sides. The discussion primarily revolved around P18’s focus on group-level legitimate factors for fairness versus P10’s approach of directly excluding the impact of protected attributes for AI decision-making fairness, consulting FEE as needed to support their arguments. P10 proposed fairness as follows: ”It’s like oh no matter the value of age, gender, foreign worker is changed or not, AI prediction should not be changed […] You eliminate the bias first and then you think about employment [and other individual qualifications]”. After this perspective was accepted, Team 3, different from Team 1 and Team 2, advocated for using Counterfactual Fairness for individual fairness first, and then using Conditional Statistical Parity with a focus on subgroup fairness. Additionally, there was unanimous agreement on the importance of including at least age and gender to secure subgroup fairness in the Conditional Statistical Parity application.
@@
Main finding 3: Teams used different strategies to come to a consensus. Voting was frequently employed but there were differences in how to decide the outcomes of voting. Reasons for or against metrics were discussed in all teams but this could also lead to entrenchments of opinions if not carefully handled. All teams chose to overcome conflict by combining metrics in some way, ending up with compromise solutions. The consensus showed that all teams considered both subgroup fairness and individual fairness, but with different priorities.